OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (1, 2)?

OpenStudy (jdoe0001):

\(\large \begin{cases} (y-{\color{blue}{ k}})^2=4p(x-{\color{brown}{ h}}) \\ (x-{\color{brown}{ h}})^2=4p(y-{\color{blue}{ k}})\\ y=(x-{\color{brown}{ h}})^2+{\color{blue}{ k}}\\ x=(y-{\color{blue}{ k}})^2+{\color{brown}{ h}} \end{cases}\qquad vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\)

OpenStudy (anonymous):

y = (x - 1) 2 - 2 is this the answer